OpenStudy (anonymous):

A ship traveling east at 25 mph is 10 mi from a harbor when another ship leaves the harbor traveling east at 35mph. How long does it take the second ship to catch up to the first ship?

OpenStudy (anonymous):

I have the answer in the back of the book, but I don't know how they came to the answer and I want to know how and not just the answer.

OpenStudy (anonymous):

ok for some reason I was thinking that you could no add the 10

OpenStudy (cwrw238):

oopss - i must be tired - i made a mistake...

OpenStudy (anonymous):

it takes the second ship 1 hour to catch the first ship is that what you got?

OpenStudy (anonymous):

but I want to know how the book got that.

OpenStudy (anonymous):

i think i got it !

OpenStudy (cwrw238):

ok the second ship will travel 10 + x miles in same time that first ship travels x miles first ship: t = x/25 hours 2nd ship t = (x + 10)/35 hours right?

OpenStudy (anonymous):

yeah

OpenStudy (cwrw238):

so x/25 = (x+ 10)/ 35 35x = 25x + 250 x = 25 t = x/25 = 25/25 = 1

OpenStudy (anonymous):

i got it !

OpenStudy (anonymous):

25(.4 + t) = 35t

OpenStudy (cwrw238):

phew - i don't know why but i struggle with these speed / time /dist problems

OpenStudy (anonymous):

i found the time of the of the first then set the distances equal

OpenStudy (cwrw238):

yes - thats a good method

OpenStudy (anonymous):

well thank you for your help! happy calculating!

OpenStudy (cwrw238):

lol - yw